### v1

# Representational Alignment Across Model Layers and Brain Regions with Hierarchical Optimal Transport

Abstract
Standard representational similarity methods align each layer of a network to its best match in another independently, producing asymmetric results, lacking a global alignment score, and struggling with networks of different depths. These limitations arise from ignoring global activation structure and restricting mappings to rigid one-to-one layer correspondences. We propose Hierarchical Optimal Transport (HOT), a unified framework that jointly infers soft, globally consistent layer-to-layer couplings and neuron-level transport plans. HOT allows source neurons to distribute mass across multiple target layers while minimizing total transport cost under marginal constraints. This yields both a single alignment score for the entire network comparison and a soft transport plan that naturally handles depth mismatches through mass distribution. We evaluate HOT on vision models, large language models, and human visual cortex recordings. Across all domains, HOT matches or surpasses standard pairwise matching in alignment quality. Moreover, it reveals smooth, fine-grained hierarchical correspondences: early layers map to early layers, deeper layers maintain relative positions, and depth mismatches are resolved by distributing representations across multiple layers. These structured patterns emerge naturally from global optimization without being imposed, yet are absent in greedy layer-wise methods. HOT thus enables richer, more interpretable comparisons between representations, particularly when networks differ in architecture or depth.
1 Introduction
Understanding high-dimensional neural activity is a shared challenge in neuroscience and artificial intelligence (AI). In neuroscience, comparing neural responses across individuals reveals which computations are universally shared versus idiosyncratic. In AI, comparing representations across models reveals how architectural choices, training objectives, and learning dynamics shape learned features, and helps identify principles of universality i.e. representational properties that emerge consistently across diverse network architectures and objectives. Comparing models to brains extends this logic further: while we cannot rerun biological evolution, we can simulate “evolution in silico” by training artificial networks with different constraints, inputs, and objectives. When such models converge on brain-like representations, they offer mechanistic hypotheses for why the brain may have adopted its computational strategies which is a deeply important theoretical question. These comparisons have revealed striking similarities between biological and artificial networks (Yamins et al. (2014); Eickenberg et al. (2017); Güçlü and Van Gerven (2015); Cichy et al. (2016); Khaligh-Razavi and Kriegeskorte (2014); Schrimpf et al. (2018; 2020); Storrs et al. (2021); Kell et al. (2018)), common computational motifs across diverse architectures and objectives Huh et al. (2024); Kornblith et al. (2019); Bansal et al. (2021); Dravid et al. (2023), and other universal representational dimensions Chen and Bonner (2025); Hosseini et al. (2024).
The standard approach to representational comparison is layer-wise matching: each source layer is paired with the single best-matched target layer under some similarity measure (e.g., Representational Similarity Analysis (Kriegeskorte et al., 2008), Centered Kernel Alignment (Kornblith et al., 2019), Procrustes distance (Williams et al., 2021) or linear predictivity). Despite its widespread use, this approach has fundamental limitations. It enforces rigid one-to-one correspondences that fail when networks differ in depth or when a source layer corresponds to features distributed across multiple target layers. It produces asymmetric layer mappings depending on the direction of comparison and yields no unified score for global network alignment. Most importantly, by optimizing each match independently, it ignores the global activation structure and risks overfitting to noise.
We propose Hierarchical Optimal Transport (HOT), a framework for globally consistent representational alignment. Optimal transport–based methods, such as Soft-Matching distance (Khosla and Williams, 2024; Khosla et al., 2024), have recently emerged as powerful metrics for comparing neural representations. Unlike metrics such as RSA, CKA, or linear predictivity—which are rotation-invariant and thus unable to capture similarities in neuron-level tuning—OT-based methods are rotation-sensitive. They explicitly match neurons based on their tuning profiles and, by relaxing hard permutation constraints into fractional couplings, can also handle layers of unequal size. This enables richer, more flexible neuron-level alignments than either rotation-invariant similarity metrics or strict permutation-based approaches. However, Soft Matching also remains limited to pairwise layer comparisons like other methods and does not capture global structure across networks. HOT fills this gap by operating hierarchically: it simultaneously infers soft neuron-to-neuron couplings within layers and a soft, globally consistent layer-to-layer coupling across the hierarchy. Rather than forcing each source layer to match exactly one target layer, HOT allows source layers to distribute their representational “mass” across multiple target layers while minimizing the total transport cost under marginal constraints; that is, each source layer must distribute exactly 100 of its mass across target layers (no information is lost), and the total mass each target layer receives from all source layers must sum to a balanced allocation (no target is over- or under-utilized). These conservation laws ensure a balanced alignment where every layer contributes meaningfully to the global correspondence, preventing any layer from being arbitrarily overweighted or ignored in the global matching. The result is a single network-level alignment score and a soft transport plan that naturally handles depth mismatches.
We evaluate HOT on three diverse domains: comparisons between foundation models in vision (Vision Transformers like DINOv2 and ViT-MAE), large language models of varying scales (LLaMA, Qwen), and fMRI recordings from human visual cortex across different participants. Our key contributions are:
A principled global alignment framework (theoretical): HOT jointly optimizes all layer correspondences to produce symmetric, globally consistent assignments with a single alignment score. In contrast, greedy pairwise approaches are asymmetric, can overweight certain layers while completely ignoring others, and and may spuriously treat wide layers as similar to every layer they are compared against, since their high dimensionality allows them to fit or partially overlap with many different representational subspaces, obscuring more meaningful correspondences.
Natural handling of depth mismatches (theoretical): By allowing soft, many-to-many mappings between layers, HOT aligns networks of different depths without forcing inappropriate one-to-one correspondences.
Rotation-invariant extension (theoretical): We propose an extension of HOT that incorporates additional orthogonal transformations, making the framework rotation-invariant. This ensures that correspondences can be recovered even when shared representational features are embedded in rotated subspaces, and yields consistently high-quality alignments.
Improved alignment scores (empirical): Across domains (vision models, large language models, and brain data), HOT matches or surpasses standard pairwise methods, yielding higher alignment.
Emergent hierarchical structure (empirical): Without imposing ordering constraints, HOT recovers known hierarchical organization in visual cortex data across subjects and reveals clean layer-to-layer correspondences in model–model comparisons where early layers map to early layers and deeper layers preserve their relative ordering. By contrast, greedy pairwise methods fail to reveal hierarchical structure, often leaving many layers unmatched while a single or few layers dominate the mappings.
Featural distribution across depth (empirical): HOT reveals how deeper networks spread computations across multiple layers that shallower networks compress into fewer stages. Concretely, a single layer in a shallower network often distributes its mass across several neighboring layers in a deeper network, an effect that greedy pairwise methods completely miss.
2 Methods
2.1 Problem Setup and Existing Approaches
Comparing the internal representations of neural networks often proceeds by a pairwise layer search under some similarity or distance measure. In general, let denote the number of stimuli (e.g., images, text sequences) used to probe both models. We consider two neural networks:
The first network has layers. Layer contains units, and its activations across the stimuli are represented by
The second network has layers. Layer contains units, with activations
Each row of or corresponds to the response of all units in that layer to a single stimulus, while each column corresponds to the activity of a single unit across stimuli. For any chosen alignment metric (e.g., linear predictivity, Procrustes distance, RSA, CKA, or Soft Matching), one typically does:
and then reports the layer-wise score . This enforces a hard one-to-one mapping from each source layer to a single target layer . However, when , or when the set of features represented in one layer of network A is distributed across multiple layers of network B, such a rigid per-layer pairing is not well-suited: a source layer may genuinely correspond to a mixture of multiple target layers. Moreover, by optimizing each layer independently, this approach ignores the global structure of all activations and can overfit to noise in any single layer’s responses.
In what follows, we introduce a hierarchical optimal transport framework that overcomes both issues by simultaneously inferring a soft layer-to-layer coupling and per-neuron transport plans, yielding a globally balanced alignment across networks of arbitrary depths (Figure 1).
2.2 Hierarchical Optimal Transport
Our framework operates at two hierarchical levels: an inner level that aligns neurons within each layer pair, and an outer level that determines how layers should be coupled globally.
2.2.1 Inner Level: Neuron-to-Neuron Transport
For each pair of layers , we compute a cost matrix
where is the correlation distance and denotes Pearson correlation. We then solve an optimal transport problem to find a coupling that minimizes the transport cost between neurons:
where is the transportation polytope:
This constraint ensures that each source neuron distributes exactly of its mass (summing rows), and each target neuron receives exactly of the total mass (summing columns), with all entries non-negative. The coupling specifies how neurons in layer can be combined to reconstruct neurons in layer .
This inner-level optimization is precisely the Soft Matching distance (Khosla and Williams, 2024) applied between layers and : it is rotation-sensitive yet permutation-invariant, and directly measures single-neuron tuning alignment. When the layers have equal size (), the optimal solution lies at a vertex of the transportation polytope; which is a permutation matrix by the Birkhoff–von Neumann theorem (von Neumann, 1953; De Loera and Kim, 2013). However, when layer sizes differ, soft matching assigns each neuron in one layer to weighted combinations of neurons in the other via soft-assignments.
2.2.2 Outer Level: Layer-to-Layer Transport
The inner costs form a layer-to-layer cost matrix . We solve a second optimal transport problem to find the global layer coupling:
where is the transportation polytope for layers:
The element represents the fraction of layer ’s representation that is explained by layer . These marginal constraints ensure mass conservation: each source layer distributes its unit mass across target layers (after normalization by ), and the total mass received by all target layers is balanced.
When the two networks have the same number of layers (), we can show that the optimal solution is a (scaled) permutation matrix: the objective is linear in , and the constraints define the transportation polytope whose vertices are permutation matrices (scaled by ) by the Birkhoff–von Neumann theorem. Since linear programs achieve their optima at vertices, the solution finds a one-to-one layer matching when . However, when , the soft coupling allows source layers to distribute their mass across multiple target layers, naturally handling depth mismatches.
2.2.3 Reconstruction and Evaluation
Given the layer coupling and neuron couplings , we reconstruct layer as:
The factor appears because sums to across . We evaluate alignment quality using the mean correlation between original and reconstructed neurons on held-out data:
The global HOT score is the average across all layers:
2.3 Rotation-Invariant Extension
Most existing metrics of representational similarity, such as RSA, CKA, and Procrustes distance, are designed to be rotation-invariant. The rationale is that when comparing population codes, we often care less about the tuning of individual neurons and more about the geometry of the representational space: distances, angles, and relative positions between stimulus responses. Two networks can encode essentially the same geometry while using different coordinate bases (for example, rotated versions of one another). Requiring strict unit-to-unit correspondence in such cases would artificially inflate dissimilarity, even though the representational geometry and information content are preserved. A rotation-invariant extension of HOT (HOT + R henceforth) therefore provides a way to capture this geometric equivalence while still enforcing a globally consistent layer- and neuron-level coupling. We introduce rotation matrices for each layer pair and minimize the reconstruction error:
We optimize via alternating minimization:
Fix , update : Solve optimal transport using correlation distance on rotated features .
Fix , update : Solve orthogonal Procrustes via SVD of .
Update : Refresh the outer coupling using the Frobenius reconstruction costs.
Predictions incorporate the learned rotations:
2.4 Baseline Comparisons
We compare HOT against several baselines to isolate the contribution of each component:
Random Layer Assignment (Perm-P):
We randomly permute the rows of , breaking the optimized layer correspondences while preserving the neuron-level optimal transport within each layer pair. This control is expected to perform reasonably well because it still finds optimal neuron matches for each (now randomly assigned) layer pairing—it only disrupts which layers are matched, not how well neurons align within those matches.
HOT Top-1 Layer Transport Plan (Single-Best OT):
Each source layer maps only to its highest-weight target layer
converting the soft layer coupling to a hard one-to-one assignment while keeping the soft neuron-level transport.
Independent Pairwise OT (Pairwise Best OT):
The standard greedy approach where each source layer is matched to the single target layer with minimum inner OT cost, computed independently on training data. Formally,
and layer is reconstructed using only . This baseline ignores global structure and can result in multiple source layers matching to the same target layer while leaving others unmatched.
Rotation-aware variants:
We evaluate rotation-invariant versions of the pairwise OT baseline (henceforth, ‘Pairwise Best + R’), where predictions use with orthogonal transformations optimized via Procrustes alignment.
3 Results
We evaluate representational similarity under the HOT metric across four distinct alignment setups: (i) large language models of different families and scales, (ii) fMRI responses from the visual cortex of four human subjects, (iii) pretrained transformer-based vision models spanning different families and scales, and (iv) cross-domain comparisons between human visual cortex and vision transformers (Appendix Section D). Across all settings, HOT matches or surpasses baseline reconstruction (prediction) scores. Importantly, the transport plans inferred by HOT naturally reveal systematic layer-wise correspondences across models and cortex, despite not being explicitly optimized for such structure; specifically, earlier layers or regions tend to align with earlier counterparts, while deeper layers or regions map to progressively higher levels; patterns that greedy pairwise methods fail to capture.
3.1 Representational Similarity Between Large Language Models
Experimental Setup. We extract layer-wise representations by averaging token activations across 2,552 prompts from the STSB dataset (May, 2021; Enevoldsen et al., 2025; Muennighoff et al., 2022). For each model, this yields a sequence of representation matrices and , corresponding to successive layers. We evaluate models of varying sizes from the LLaMA-3.2 (Grattafiori et al., 2024) and Qwen-2.5 (Qwen et al., 2025) families. Representations are compared using HOT and baseline metrics, and the resulting transport plans are analyzed. To quantify alignment quality, we reconstruct representations on a held-out validation split (20) using the learned transport maps and report the correlation with ground-truth activations, as detailed in the Methods Section2.2.3.
| Model 1 | Model 2 | HOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
| Llama-3.2 1B | Llama-3.2 3B | 0.558 | |||
| Qwen-2.5 0.5B | Qwen-2.5 3B | 0.510 | |||
| Qwen-2.5 0.5B | Llama-3.2 1B | 0.522 | |||
| Qwen-2.5 0.5B | Llama-3.2 3B | 0.531 | |||
| Llama-3.2 1B | Qwen-2.5 3B | 0.432 | |||
| Llama-3.2 3B | Qwen-2.5 3B | 0.383 |
Results. HOT consistently achieves higher reconstruction accuracy than baseline methods. As shown in Table 1, reconstruction scores on the held-out validation set are significantly higher under HOT, indicating that the alignment plans it learns are more robust and generalizable than those obtained from pairwise baselines. This improvement stems from HOT’s (i) enforcement of global consistency across all layers and (ii) ability to distribute representational mass across multiple target layers, allowing it to recover alignments that remain hidden to pairwise methods when features from a single layer are distributed across several layers in another model. Beyond quantitative gains, HOT uncovers clear hierarchical correspondences between models. As illustrated in Figures 2 and A.1, the transport plans produced by HOT exhibit strong diagonal structure: early layers in one model align with early layers in the other, while deeper layers align with deeper layers. Such structured correspondences are absent in pairwise OT, which often yields noisy mappings. Furthermore, when comparing shallower with deeper models, HOT reveals that single layers in the shallower model distribute their mass across multiple consecutive layers in the deeper model. This soft many-to-many mapping reflects how additional depth refines and spreads computations, suggesting that representational stages compressed into one layer in a shallower model are decomposed across multiple processing steps in a deeper model—an organizational principle that greedy baselines fail to uncover.
3.2 Representational Similarity Between Visual Cortex across Subjects
Experimental Setup. We analyze fMRI responses from the Natural Scenes Dataset (NSD; (Allen et al., 2022)), which contains recordings from 8 participants who each viewed up to 10,000 natural images. Of these, 4 subjects viewed the full set of 10,000 images three times, with 1,000 images shared across all participants. We focus on these 4 subjects and restrict our analysis to their responses to the shared 1,000 images, ensuring that alignment is evaluated on common stimuli.
We target visual cortex regions (V1–V4), treating each region as a “layer” and individual voxels as “neurons”. The visual cortex provides a strong testbed for hierarchical alignment because it is one of the best-characterized cortical systems: early areas (V1, V2) are known to encode low-level features such as orientation and contrast, while higher areas (V3, V4) encode progressively more complex shapes and object features (Hubel and Wiesel, 1968; Pasupathy and Connor, 2001; Desimone and Schein, 1987; Desimone et al., 1984). This hierarchical progression is well-established across individuals, so correspondences between homologous regions are strongly expected. HOT is applied to align responses across subjects by generating transport maps between regions. To evaluate alignment quality, we reconstruct held-out responses (20 validation split) and report correlations with ground-truth activity. We repeat this evaluation across 5 random train-validation splits to ensure robustness.
| Model 1 | Model 2 | HOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
| Subject A | Subject B | 0.308 | |||
| Subject A | Subject C | 0.232 | |||
| Subject A | Subject D | 0.253 | |||
| Subject B | Subject C | 0.237 | 0.237 | 0.237 | |
| Subject C | Subject D | 0.239 | |||
| Subject B | Subject D | 0.254 |
Results. As shown in Tables 2 and B.1, HOT achieves reconstruction scores comparable to pairwise OT, with only a marginal decrease in correlation. More importantly, Figures 3 and B.1 show that the transport maps inferred by HOT recover the expected cross-subject correspondences: cortical regions in one subject consistently align to the same regions in another. In contrast, pairwise OT does not produce such structured mappings in any subject pair. This indicates that HOT’s global alignment scheme captures region-to-region correspondences that pairwise methods miss. To rule out the possibility that this effect stems from the OT optimization framework itself rather than from the layer-wise setup, we additionally perform pairwise linear predictivity optimization. As shown in Figure B.2, the best-matching layers identified by linear predictivity also do not align corresponding regions, further confirming that the hierarchical alignment mechanism in HOT is critical for recovering robust and generalizable correspondences. Overall, these results demonstrate that HOT yields transport plans that are both interpretable and biologically meaningful, while maintaining reconstruction performance on par with baseline methods.
3.3 Representation Similarity between Vision Models
Experimental Setup. We extract layer-wise representations from Vision Transformers by averaging patch activations for each input image. We use 20,000 images randomly sampled from the ImageNet validation set (Deng et al., 2009; Russakovsky et al., 2015), sampled to ensure a uniform coverage across all classes in Imagenet. For each model, this yields a sequence of representation matrices that serve as inputs to the HOT framework. We evaluate two families of pretrained vision transformers—DINOv2 and ViT-MAE across multiple model scales. Alignment quality is quantified by reconstructing representations on a held-out validation split (20) using the learned transport plans, with reconstruction–ground truth correlation as the metric.
Prior work has shown that the residual stream in Transformers lacks privileged axes and is invariant up to rotations of its basis (Khosla et al., 2024). Since HOT, like other OT-based methods, is rotation-sensitive, we additionally evaluate a rotation-augmented variant (HOT+R) and its baselines (see Methods Section 2.3). We restrict this analysis to Vision Transformers, where rotational invariances are especially relevant and where the computational cost of HOT+R remains tractable. For large language models and fMRI data, the added optimization over rotations was computationally prohibitive, so we report only the rotation-sensitive results in those domains. In HOT+R, the learned rotation matrices are incorporated into both transport optimization and evaluation, ensuring that geometric equivalences induced by rotations are properly captured.
| Model 1 | Model 2 | HOT Metric | Pairwise Best OT | HOT + R | Pairwise Best + R |
| DINOv2 Small | ViT-MAE Base | 0.600 | |||
| DINOv2 Small | DINOv2 Large | 0.778 | |||
| DINOv2 Small | DINOv2 Giant | 0.790 | |||
| DINOv2 Small | ViT-MAE Large | 0.633 | |||
| DINOv2 Small | ViT-MAE Huge | 0.657 | |||
| ViT-MAE Base | DINOv2 Large | 0.732 | |||
| ViT-MAE Base | DINOv2 Giant | 0.580 | |||
| ViT-MAE Base | ViT-MAE Large | 0.850 | |||
| ViT-MAE Base | ViT-MAE Huge | 0.788 | |||
| ViT-MAE Huge | DINOv2 Giant | 0.614 |
Results. Table 3 shows that, on its own, vanilla HOT does not consistently outperform pairwise OT in reconstruction accuracy. Consistent with this, the transport plans inferred by HOT (Figure 4 and Figures C.1 - C.10) only partially reveal layer-wise correspondences: in some cases, clear diagonal structure emerges, but in others the mappings are noisier. By contrast, the rotation-augmented variant (HOT+R) yields substantially higher reconstruction scores surpassing both vanilla HOT and pairwise vanilla and rotational baselines (Tables 3, C.1 and C.2). Importantly, the transport plans produced by HOT+R consistently exhibit strong hierarchical correspondences, recovering clean layer-to-layer alignment even in settings where vanilla HOT fails to do so. These findings indicate that incorporating rotation into the OT framework not only improves quantitative alignment quality but also produces more generalizable and interpretable mappings, particularly in domains like Vision Transformers where representations are known to be rotation-invariant.
4 Discussion
Hierarchical OT provides a principled framework for aligning two networks with arbitrary depths. By softly coupling all layer objectives, HOT (i) allows a neuron in one layer of a source network to align with a soft combination of units distributed across multiple layers of a target network, (ii) enforces global consistency that mitigates overfitting to noise in any single layer, and (iii) produces a single network-level—rather than layer-level—alignment score. In our experiments, HOT achieves higher alignment scores than greedy pairwise matches and yields intuitive transport plans that reveal hierarchical correspondences both in artificial networks and in the human brain.
Limitations.
Despite these strengths, several limitations remain. First, HOT is computationally demanding: solving the inner OT scales as for neurons in a pair of layers, and the overall procedure is for networks with layers. This makes scaling to very wide or deep models challenging without further algorithmic improvements. Second, our evaluation was limited to a subset of models and brain datasets; follow-up analyses on more models and diverse neural data will be critical to assess generality. Finally, while HOT provides a descriptive alignment, it does not by itself explain why certain features converge across systems, or which computational principles drive these correspondences.
Future directions.
Several extensions are natural. One avenue is to add a third hierarchical level to capture representational dynamics over the course of training, enabling alignment not only across neurons and layers but also across developmental or learning trajectories. Applications could include aligning models across different stages of training, or comparing biological learning to gradient descent. Another is to incorporate priors on the transport plan (e.g. smoothness) to guide alignment toward more interpretable solutions. Ultimately, as both biological and artificial networks grow in scale and complexity, methods like HOT that respect global structure while revealing network correspondences will be essential for understanding the universal principles governing intelligent systems.
Appendix
Appendix A Comparing LLM-LLM Representations
Appendix B Representation Similarity between Vision Cortex Response
| Model 1 | Model 2 | HOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
| Subject A | Subject B | 0.306 0.007 | 0.258 0.008 | 0.306 0.007 | 0.308 0.007 |
| Subject A | Subject C | 0.227 0.008 | 0.184 0.008 | 0.227 0.008 | 0.232 0.009 |
| Subject A | Subject D | 0.248 0.014 | 0.212 0.016 | 0.248 0.014 | 0.253 0.014 |
| Subject B | Subject C | 0.237 0.008 | 0.199 0.006 | 0.237 0.008 | 0.237 0.009 |
| Subject C | Subject D | 0.235 0.010 | 0.196 0.009 | 0.235 0.010 | 0.239 0.008 |
| Subject B | Subject D | 0.248 0.013 | 0.217 0.011 | 0.248 0.013 | 0.254 0.012 |
B.1 Transport plans for OT based mapping pairs
B.2 Transport plans for Linear Predictivity based mapping pairs
Appendix C Representation Similarity between Vision Models
| Model 1 | Model 2 | HOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
| DINOv2 Small | ViT-MAE Base | 0.301 | |||
| DINOv2 Small | DINOv2 Large | 0.353 | |||
| DINOv2 Small | DINOv2 Giant | 0.466 | |||
| DINOv2 Small | ViT-MAE Large | 0.381 | |||
| DINOv2 Small | ViT-MAE Huge | 0.411 | |||
| ViT-MAE Base | DINOv2 Large | 0.624 | |||
| ViT-MAE Base | DINOv2 Giant | 0.202 | |||
| ViT-MAE Base | ViT-MAE Large | 0.598 | |||
| ViT-MAE Base | ViT-MAE Huge | 0.417 | |||
| ViT-MAE Huge | DINOv2 Giant | 0.352 |
| Model 1 | Model 2 | HOT + R | Single-Best + R | Pairwise Best + R |
| DINOv2 Small | ViT-MAE Base | 0.600 | 0.600 | |
| DINOv2 Small | DINOv2 Large | 0.778 | ||
| DINOv2 Small | DINOv2 Giant | 0.790 | ||
| DINOv2 Small | ViT-MAE Large | 0.633 | ||
| DINOv2 Small | ViT-MAE Huge | 0.657 | ||
| ViT-MAE Base | DINOv2 Large | 0.732 | ||
| ViT-MAE Base | DINOv2 Giant | 0.580 | ||
| ViT-MAE Base | ViT-MAE Large | 0.850 | ||
| ViT-MAE Base | ViT-MAE Huge | 0.788 | ||
| ViT-MAE Huge | DINOv2 Giant | 0.614 |
Appendix D Representation Similarity between Vision Models and Vision Cortex
Experimental Setup. We assess representational similarity between human visual cortex and vision transformers by comparing fMRI responses with model activations elicited by the same set of stimuli. Specifically, we use responses from the Natural Scenes Dataset ([Allen et al., 2022]), focusing on 1,000 shared images viewed by participants in the fMRI experiment. The same images are presented to pretrained vision transformers, and layer-wise representations are extracted by averaging patch embeddings across each input. Following the approach used in the cortex–cortex analysis, we treat distinct visual areas (V1–V4) as “layers” and individual voxels as “neurons.” We then compute both HOT and its rotation-augmented variant (HOT+R) to align cortical responses with model representations. This design enables a direct comparison of hierarchical organization across biological and artificial systems under matched visual input.
| Model 1 | Model 2 | HOT Metric | Pairwise Best OT | HOT + R | Pairwise Best + R |
| Subject | DINOv2 Base | 0.145 | |||
| Subject | DINOv2 Giant | 0.163 | |||
| Subject | ViT-MAE Base | 0.151 | |||
| Subject | ViT-MAE Huge | 0.154 |
Results. Table D.1 reports reconstruction scores for HOT, HOT+R, and corresponding baselines averaged across the four subjects. Among all methods, HOT+R achieves the highest reconstruction score, indicating that incorporating rotation is critical for capturing shared structure between cortical and model representations. Subject-specific results (Tables D.2 and D.3) further confirm this trend, with HOT+R consistently outperforming both vanilla HOT and pairwise baselines. Transport plans visualized in Figures D.1– D.16 reveal partial but inconsistent layer-wise correspondences: in some cases, early cortical regions align with early model layers and higher regions map to deeper layers, while in others the mappings appear noisier.
| Model 1 | Model 2 | HOT Metric | Random (Perm-P) | Single-Best OT | Pairwise Best OT |
| Subject A | DINOv2 Giant | ||||
| Subject A | DINOv2 Base | ||||
| Subject A | ViT-MAE Huge | ||||
| Subject A | ViT-MAE Base | ||||
| Subject B | DINOv2 Giant | ||||
| Subject B | DINOv2 Base | ||||
| Subject B | ViT-MAE Huge | ||||
| Subject B | ViT-MAE Base | ||||
| Subject C | DINOv2 Giant | ||||
| Subject C | DINOv2 Base | ||||
| Subject C | ViT-MAE Huge | ||||
| Subject C | ViT-MAE Base | ||||
| Subject D | DINOv2 Giant | ||||
| Subject D | DINOv2 Base | ||||
| Subject D | ViT-MAE Huge | ||||
| Subject D | ViT-MAE Base |
| Model 1 | Model 2 | HOT + R | Single-Best + R | Pairwise Best + R |
| Subject A | DINOv2 Giant | |||
| Subject A | DINOv2 Base | |||
| Subject A | ViT-MAE Huge | |||
| Subject A | ViT-MAE Base | |||
| Subject B | DINOv2 Giant | |||
| Subject B | DINOv2 Base | |||
| Subject B | ViT-MAE Huge | |||
| Subject B | ViT-MAE Base | |||
| Subject C | DINOv2 Giant | |||
| Subject C | DINOv2 Base | |||
| Subject C | ViT-MAE Huge | |||
| Subject C | ViT-MAE Base | |||
| Subject D | DINOv2 Giant | |||
| Subject D | DINOv2 Base | |||
| Subject D | ViT-MAE Huge | |||
| Subject D | ViT-MAE Base |
Appendix E Use of Large Language Models
LLMs were used in this work to assist with writing tasks, specifically for ensuring grammatical correctness and enhancing the clarity of the text.